# Sharp Bounds for Multiple Models in Matrix Completion

### v2

@@ 4.8 Other technical lemmas
$\|\bar{A}\|_{1}$ $=\|P_{A_{0}}(\bar{A}-A_{0})+P_{A_{0}}^{\perp}(\bar{A}-A_{0})+P_{A_{0}}(A_{0})\|_{1}$
$\geq\|P_{A_{0}}^{\perp}(\bar{A}-A_{0})\|_{1}+\|P_{A_{0}}(A_{0})\|_{1}-\|P_{A_{0}}(\bar{A}-A_{0})\|_{1}.$
Then, we get
$0\leq\frac{4}{3}\|P_{A_{0}}(\bar{A}-A_{0})\|_{1}-\frac{2}{3}\|P_{A_{0}}^{\perp}(\bar{A}-A_{0})\|_{1}.$
For (2), $L(A)=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}$ and $\nabla L(A_{0})=\frac{2}{n}\sum_{i=1}^{n}\xi_{i}X_{i}$. For (8), $L(A)=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\langle X_{i},A\rangle)^{2}}$ and $\nabla L(A_{0})=\frac{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}}{L(A_{0})}=\frac{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}X_{i}}{\sqrt{\frac{1}{n}\sum_{i=1}^{n}\xi_{i}^{2}}}$. For (10), $L(A)=\frac{1}{n}\sum_{i=1}^{n}l_{\tau}(Y_{i}-\langle X_{i},A\rangle)$, and $\nabla L(A_{0})=\frac{1}{n}\sum_{i=1}^{n}\psi_{\tau}(\xi_{i})X_{i}$. ∎

### v3

@@ 4.6 Concentration of Maximum of Independent Random Variables
$\sup_{i\in\{1,\cdots,n\}}|\xi_{i}|\leq|\xi_{1}|+\sup_{i\in\{1,\cdots,n\}}|\xi_{i}-\xi_{1}|.$
By (34) we get
$\mathbb{P}\left(\sup_{i\in\{1,\cdots,n\}}|\xi_{i}|\geq e^{1/\alpha}\left(C_{1,\alpha}\sigma(\log n)^{1/\alpha}+(C_{2,\alpha}+1)\sigma t\right)\right)$
$\leq$ $\mathbb{P}\left(\sup_{i\in\{1,\cdots,n\}}|\xi_{i}-\xi_{1}|\geq e^{1/\alpha}\left(C_{1,\alpha}\sigma(\log n)^{1/\alpha}+C_{2,\alpha}\sigma t\right)\right)+\mathbb{P}\left(|\xi_{1}|\geq e^{1/\alpha}\sigma t\right)$
$\leq$ $\exp(-t^{\alpha}/\alpha)+\frac{2}{e}\exp(-t^{\alpha}/\alpha)\leq 2\exp(-t^{\alpha}/\alpha),$
where we use $\mathbb{P}(X+Y\geq a+b)\leq\mathbb{P}(X\geq a)+\mathbb{P}(Y\geq b)$. In a cleaner form, we present the concentration inequality
$\mathbb{P}\left(\sup_{i\in\{1,\cdots,n\}}|\xi_{i}|\geq C_{1,\alpha}\sigma(\log n)^{1/\alpha}+C_{2,\alpha}\sigma t^{1/\alpha}\right)\leq 2\exp(-t)$ (35)
As a consequence, if $\{\xi_{i}\}_{i=1}^{n}$ is a sequence of sub-exponential random variables satisfying $\|\xi_{i}\|_{\psi_{2}}\leq\sigma$, there exist constants $C_{1},C_{2}$ such that
$\mathbb{P}\left(\sup_{i\in\{1,\ldots,n\}}|\xi_{i}|\geq C_{1}\sigma\sqrt{\log n}+C_{2}\sigma\sqrt{t}\right)\leq 2\exp(-t).$ (36)
Lemma 16.
Let $Y$ be a random variable with mean 0 and $\mathbb{E}[Y^{2}]\leq\sigma^{2}$. For any $y$,
$\operatorname{Var}(\phi_{y}(Y))\leq\sigma^{2}.$ (37)
Proof.
Let $Y^{\prime}$ be an independent copy of $Y$. Since the function $\psi_{y}(\cdot)$ is Lipchitz with Lipchitz constant $1$,
$2\operatorname{Var}(\psi_{y}(Y))\leq\mathbb{E}(\psi_{y}(Y)-\psi_{y}(Y^{\prime}))^{2}\leq\mathbb{E}(Y^{\prime}-Y)^{2}=2\mathbb{E}[Y^{2}].$
Then, we conclude the lemma. ∎